Mauro Products distributes a single product, a woven basket whose selling price is $14 per unit and whose variable expense is $10 per unit. The company’s monthly fixed expense is $12,000.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

1. Break-even point in unit sales = $ 3,000 per unit

2. Break-even point in dollar sales = $ 42,002.10

3. New break-even point in unit sales = $ 3,150 per unit

New Break-even point in dollar sales = $ 44,102.21

Explanation

1. Contribution per unit = Selling price – Variable expense per unit

= 14 - 10

= $4 per unit

  Break-even point in unit sales = Fixed expense / Contribution per unit

= $12,000 / $4 per unit

= $ 3,000 per unit

2. Contribution margin ratio = Contribution per unit / Selling price

= 4 / 14

= 28.57%

  Break-even point in dollar sales = Fixed expense / Contribution margin ratio

= $12,000 /  28.57%

= $ 42,002.10

3. If the company’s Fixed expenses increase by $600

New Break-even point in unit sales = Fixed expense / Contribution per unit

= (12,000 + 600) / 4

= $ 3,150 per unit

New Break-even point in dollar sales = Fixed expense / Contribution margin ratio

= (12,000 + 600) / 28.57%

= $ 44,102.21